OpenStudy (anonymous):

Suppose F(xy)=x^2i+y^2j and C is the line segment segment from point P=(1,5) to Q=(−3,10). I got the vector parametric equation as r(t)=<1-4t,5+5t>. I just dont understand this part:Using the parametrization in part (a), the line integral of F along C is intergal from (a to b) of F(r(t)) dot product of r prime(t). I know the limits of integration are a=0 and b=1. I just dont know how to write the integral.

OpenStudy (anonymous):

\[ x(t)=1-4t\\ y(t) = 5 +5(t)\\ r'(t)=(x'(t),y'(t))=(-4,5)\\ F(x(t),y(t))=\left((1-4 t)^2,(5 t+5)^2\right)\\ \] Compute the dot product of F(x(t),y(t)) and r'(t) and you integrate what you obtain from a to b with respect to t

OpenStudy (anonymous):

\[ F(x(t),y(t)).r'(t)=5 (5 t+5)^2-4 (1-4 t)^2=61 t^2+282 t+121 \]

OpenStudy (anonymous):

Your final step should be \[ \int_0^1 \left(t^2+282 t+121\right) \, dt=\frac{787}{3} \]